Operations Report – December 9th

Crew 200 Operations Report 09-Dec-2018

SOL: 8

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 6:15 pm last night to 9:20 am. On tonight at 5:15 pm.

Solar— SOC 84%
Diesel Reading – 25 %
Propane Reading – 35 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 400 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – refilled, hgih level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 3

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours:
Beginning charge:
Ending charge: 73%
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: Rovers were not used this week
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? Yes, to go to Hanksville at 3:00 pm to refill water, buy gas for the car and throw our garbage
CrewCar used and why, where? No

General notes and comments: We are restarting sim (as part of our extended 3-member crew simulation) this evening at 7:00 pm.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: we filled the water tank and put some more tarps on the tunnels with Shannon.

Summary of GreenHab operations: We filled the water tank

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report [end]

Greenhab Report – December 9th

Crew 200 Greenhab Report 09-DEC-2018

Greenhab Officer:
Makiah Eustice

Environmental Control:
Heating
40% Shade Cloth on

Average Temperatures:
Low: 20 C
High: 54 C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
6 gallons

Water in Blue Tank (lbs):
70%

Times of Watering for Crops:
1100, 1830

Changes to crops:
NONE

Narrative:
We refilled the water tank. It’s pretty full now.

Harvest:
NONE

Support/Supplies Needed:
Any ideas for sensors to put in the water tank to know the level? [end]

Journalist Report – December 9th

Crew 200 Journalist Report 09-DEC-2018

I awoke to see light through the cracks of my door, the sun shining and sweeping the living room with hollow brightness. Three chairs wait our for rising, sleepy astronauts. The coffee pot and tea kettle sit for the order of 1 green tea, 2 black coffees, no more. The water tank pump has a moment of relief with the reduced crew. And saying the toilet was in better shape is an understatement.

The astronauts awoke, and the hollowness filled with warmth.

Antoine, Lindsay, and I enjoy a casual morning, discussing the evolving commercial space industry and our new roles. We got lost in conversation with the MDRS Director about the rich history of this place we can temporarily call home. I look to my crewmates, now peers, and see the same inspiration and dedication to make this place better. This spirit carried through the rest of the day, as we collected water for the static tank, dumped garbage, and helped repair the tunnel.
As we plan to do the first EVA ever done with a 3 member crew. As we do the jobs of seven.
We are young, fresh in our understanding of Mars, the space industry, and leadership. But we are adaptable, ready for the challenges of today, and dedicated to the future of innovation across borders and cultures. We are truly the generation that will carry humanity to the red planet.

We are ready to awake the Earth when we return.

Makiah Eustice

Sol Summary – December 9th

Crew 200 Sol Summary Report 9-Dec-2018

Sol: 7

Summary Title: Reduced crew EVA planning, replenishing supplies, and fixing tunnels

Author’s name: Lindsay Rutter

Mission Status: Nominal

Sol Activity Summary: This morning we rejuvenated a bit. We ate breakfast two hours past our usual meal time. We had a long and engaging conversation with Shannon on a variety of topics, including how to conduct EVAs during this reduced-crew mission. As today was our transition day between our nominal and extended missions, we returned to Hanskville to replenish gas and water and to throw out the non-burnable trash. After that, we fixed some of the tunnels connecting to the RAM and GreenHab to make sure they are extra safe. We can’t wait to to return to Mars this evening for our extended mission!

Look Ahead Plan: We plan to safely conduct our first reduced-crew EVA tomorrow and begin to map out the radio communication zones around the habitat. We will also keep track of what methods work best for thoroughly cleaning the ScienceDome floor. We brainstormed numerous project ideas today and will begin to focus on the most practical and meaningful ones starting tomorrow.

Anomalies in work: No anomalies.

Weather: Clear, sunny, and cool

Crew Physical Status: Nominal

EVA: N/A

Reports to be filed: N/A

Support Requested: None

Astronomy Report – December 9th

Crew Astronomer: Andrew Foster

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: None.

Requested objects – None.

Objects Viewed: None

Images attached:

1. Dark sky from MDRS: Sirius, Orion, Taurus and winter Milky Way. Canon D60, 40x15s images + 4 darks @ ISO5000, images processed w/ DeepSkyStacker and PS.

2. Solar feature, likely faint sunspot AR2729, captured with Lunt 100mm solar scope 08Dec2018, w/ Imaging Source DBK colour camera @15fps, best 1200 of 1400 frames, processed w/ Registax 6 & PS.

Problems Encountered:

N/A

MUSK OBSERVATORY

Solar Features Observed (08DEC2018): Solar filament, faint Sunspot AR2729. Note – found observation of solar features challenging, likely due to inexperience with coarse / fine tuning of solar etalon.

Notes

Crew Astronomer departed station 12-08-2018, (now post-processing images acquired during mission).

End report.